OpenStudy (marcelie):

find a taylor series of f(x)= cos(x) at x=pi/3

OpenStudy (marcelie):

@zepdrix

OpenStudy (marcelie):

@Directrix

OpenStudy (marcelie):

i got to here f(x) = cos (x) = 1/2 f'(x) = -sin(x) = - rad 3 /2 f''(x) = cos (x) = -1/2 f'''(x) = sin (x) = rad 3 /2

OpenStudy (marcelie):

come back D: @IrishBoy123

OpenStudy (marcelie):

oh so then how should it be ?

OpenStudy (marcelie):

ik there is formula first but we find 3 derivatives of this right?

OpenStudy (marcelie):

please help ...@steve816

OpenStudy (kmeis002):

So far your derivatives are fine. Can we start to define a rule for \( \frac{d^nf}{dx^n}\)?